Part A How many oscillations will this pendulum make before its amplitude has decreased to 15% ofliitial ampitude? Express your answer as an integer แ. Review 1 Constants Periodic Table A damped pendulum has a period of 0 54 s and a time constant of 3.8 B
Answer #1

as per damping phenomenon

percentage = e ^( - t / T)

0.15 = e^(- t/ 3.8)

t = 7.21 second

number of oscillation

n = 7.21 / 0.54

n = 13.35≈ 13

======